Marketing estimates that a new instrument for the analysis of soil samples will be very successful, moderately successful, or unsuccessful, which is probabilities 0.3, 0.6, and 0.1, respectively. The yearly revenue associated with a very successful, moderately successful, or unsuccessful is $10 million, $5 million, and $1 million, respectively. Let the random variable X denote the yearly revenue of the product. Determine:




a) the mass function of X;




b) the mean of X; and




c) the variance of X.

1
Expert's answer
2022-03-16T17:05:16-0400

Solution

Let X denote the yearly revenue of the product.


(a) The mass function of X

The probability mass function of X is as shown in the diagram below.





(b) The mean of X


Mean "\mu=\sum XP(X)"


"\mu=10(0.3)+5(0.6)+1(0.1)"


"\mu=6.1"


(c) The variance of X


"\sigma^2=\sum X^2.P(X)-\mu"


"\sigma^2=[10^2(0.3)+5^2(0.6)+1^2(0.1)]-6.1^2"


"\sigma^2=45.1-37.21"


"\sigma^2=7.89"